a library expansion, began in 2002, was expected to cost $109 million. By 2006, library officials estimate the cost would be $35 million over budget. By what percent did the actual cost exceed the initial estimate?
1
Expert's answer
2012-11-27T06:58:14-0500
Solution:
over = (109+35)/109 - 1 = 1.3211-1 = 0.3211 (32.11%)
Answer:
Actual cost exceed the initial estimate by 32.11%